OpenStudy (anonymous):

Which set of integers is not a Pythagorean triple? A. 7, 24, 25 B. 13, 84, 85 C. 10, 24, 25 D. 20, 21, 29

Vocaloid (vocaloid):

are you familiar with the pythagorean theorem?

OpenStudy (danjs):

the triples make a right triangle right

OpenStudy (anonymous):

no i dont vocaloid i am not familiar with it im sorry

Vocaloid (vocaloid):

the pythagorean theorem states that, for a right triangle a^2 + b^2 = c^2

OpenStudy (anonymous):

ok...

Vocaloid (vocaloid):

so, if we want to see whether three numbers make a pythaogrean triple, we plug them into the equation and see if both sides are equal

OpenStudy (anonymous):

...

OpenStudy (danjs):

look up the definition of the triple first, prollyu

OpenStudy (anonymous):

me no understand

Vocaloid (vocaloid):

let's take A for example: 7, 24, 25 we can let a = 7, b = 24, and c = 25 7^2 + 24^2 = 25^2 625 = 625 both sides are equal, so this is a pythagorean triple is that clear?

OpenStudy (anonymous):

a lil

Vocaloid (vocaloid):

do you think you can try to test B on your own?

OpenStudy (anonymous):

its 576

OpenStudy (anonymous):

b is 576

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

13^2=169 84^2 = 7056 85=7225

Vocaloid (vocaloid):

good, now add a^2 + b^2

Vocaloid (vocaloid):

169+7056 = 7225, which is equal to c^2, so c is not the answer now try answer choice C

OpenStudy (anonymous):

10^2=100 24^2=576 25^2=625

OpenStudy (anonymous):

100+576=676 not 625 it could be c